OpenStudy (anonymous):

Which of the following conditions need to hold for two functions f and g to be equal? Check all that apply. A. The values of f and g are equal at each point in the domain. B. The domains of f and g are equal. C. The rules for f and g are equal. D. The ranges of f and g are equal.

OpenStudy (anonymous):

C D A

OpenStudy (anonymous):

nah man, its not a check off one bruh, its a mutipule choice one

OpenStudy (anonymous):

nevermind you right bru thanks

geerky42 (geerky42):

I don't know what C means, but I think B should be checked. Saying you have function \(f(x) = \dfrac{x^2-25}{x-5}\) and \(g(x) = x+5\). If you simplify f(x), you will get g(x), but they are not exact equal since \(f(x)\neq g(x)\) at \(x=5\) because \(f(x)\) doesn't exist at \(x=5\). Therefore, for function to be equal, they need to have equal domain... ... and the values of f and g to be equal at each point in the domain. Hence my answer is A B D. Not sure about C though...

OpenStudy (anonymous):

thank yall so much (':